Solve and then check. Show ALL work
2x^2+3x-24=5x

Answer by rothauserc(4718)

2x^2 +3x -24 = 5x
subtract 5x from both sides of =
2x^2 -2x -24 = 0
divide both sides of = by 2
x^2 -x -12 = 0
factor this polynomial
( x - 4 ) * ( x + 3 ) = 0
x - 4 = 0
x = 4
x + 3 = 0
x = -3
therefore,
x = 4 or -3
check answers by substituting for x in original equation
*********************************
x = 4
2x^2 +3x -24 = 5x
2*(4^2) +3*(4) -24 = 5*(4)
32 +12 - 24 = 20
44 - 24 = 20
20 = 20
x = 4 checks
***********************************
x = -3
2x^2 +3x -24 = 5x
2*(-3^2) +3*(-3) -24 = 5*(-3)
18 -9 -24 = -15
18 - 33 = -15
-15 -15
x = -3 checks